# False detection rate control in time series coincidence detection

### v1

@@ 3.2.2 A guarantee under stationarity and $\beta$-mixing
Importantly, in the result of Theorem 2, the upper bound does not vary with tt—it offers a meaningful bound even for the “real-time” detection problem, when we aim to verify that an event has occurred at time $t=T$. Of course, this comes at the cost of requiring the additional $\beta$-mixing assumption. In Theorem 1, on the other hand, since we do not assume any mixing conditions, then we need to be able to observe data occurring both before time $t$ and after time $t$ (i.e., $t$ should lie in the middle of the time range), to be able to assess whether the evidence observed at time $t$ is significant.
@@ 6 Discussion and related work
In this work, we have examined the problem of coincidence detection in time series data, in settings where accurate models for the data distribution may not be available. Our results show that time-shifting methods offer false positive control under only weak assumptions on the properties of the time series data. These findings support the existing use of time-shifting inference procedures across a range of scientific applications.
@@ Acknowledgements
We gratefully acknowledge the support of the NSF-Simons AI-Institute for the Sky (SkAI) via grants NSF AST-2421845 and Simons Foundation MPS-AI-00010513. Additionally, R.L. and R.F.B. were partially supported by the National Science Foundation via grant DMS-2023109, R.F.B. was partially supported by the Office of Naval Research via grant N00014-24-1-2544. S.D. and D.E.H. were supported by the National Science Foundation via grants PHY-2110507 and PHY-2513312. D.E.H. was also supported by the Simons Collaboration on Black Holes and Strong Gravity. The authors thank Brent Doiron, Margaret Gardel, and Elizabeth Jerison for helpful discussions.

### v2

# False positive control in time series coincidence detection

@@ 3.2.1 A guarantee under stationarity
In Appendix B, we verify that this inflation, where the probability of a false positive may be larger than the nominal level $\alpha$, cannot be avoided without further assumptions. In particular, when $\alpha$ is small, the inflation factor of this upper bound is tight for any $t$: for example, if testing at the midpoint ($t\approx T/2$), it is indeed possible to obtain a false positive probability $\approx 4\alpha$ for certain distributions on the data.
@@ 3.2.2 A guarantee under stationarity and $\beta$-mixing
Importantly, in the result of Theorem 2, the upper bound does not vary with tt—it offers a meaningful bound even for the “real-time” detection problem, when we aim to verify that an event has occurred at time $t=T$. Of course, this comes at the cost of requiring the additional $\beta$-mixing assumption, which is required to ensure that any dependence within a single data stream must be relatively short-lived. In Theorem 1, on the other hand, since we do not assume any mixing conditions, then we need to be able to observe data occurring both before time $t$ and after time $t$ (i.e., $t$ should lie in the middle of the time range), to be able to assess whether the evidence observed at time $t$ is significant.
@@ 6 Discussion and related work
In this work, we have examined the problem of coincidence detection in time series data, in settings where accurate models for the data distribution may not be available. Our results show that time-shifting methods offer false positive control under only weak assumptions on the properties of the time series data. While the guaranteed upper bounds on the false positive probability do have some inflation above the target level $\alpha$, we emphasize that in practice we would often expect to see false positive probability $\approx\alpha$ (as is indeed the case in many of our simulation settings); we can interpret our results as providing reassurance that, even in the worst case, any inflation of the false positive probability cannot be too severe. These findings support the existing use of time-shifting inference procedures across a range of scientific applications.
@@ Acknowledgements
We gratefully acknowledge the support of the NSF-Simons AI-Institute for the Sky (SkAI) via grants NSF AST-2421845 and Simons Foundation MPS-AI-00010513. Additionally, R.L. and R.F.B. were partially supported by the National Science Foundation via grant DMS-2023109, R.F.B. was partially supported by the Office of Naval Research via grant N00014-24-1-2544. S.D. and D.E.H. were supported by the National Science Foundation via grants PHY-2110507 and PHY-2513312. D.E.H. was also supported by the Simons Collaboration on Black Holes and Strong Gravity through grant SFI-MPS-BH-00012593-07. The authors thank Brent Doiron, Margaret Gardel, and Elizabeth Jerison for helpful discussions.
@@ A.5.3 Proof of Lemma 3
Appendix B An example for matching the inflation factor in Theorem 1
In the simulation experiments in Section 5, the p-values for testing the midpoint in time (i.e., $p_{t}$, for testing whether an event occurred at time $t\approx T/2$) result in a false positive probability that is approximately bounded by $\alpha$, empirically. In contrast, Theorem 1 only establishes an upper bound of $\approx 4\alpha$ on the probability of a false positive. This raises the question of whether the theoretical bound is loose. However, we will show with examples that the inflation factor $1/\mathsf{mar}(t)$ in Theorem 1 is tight (when $\alpha$ is sufficiently small). In particular, when $t\approx T/2$, the factor of $4$ cannot be improved. For simplicity, we will work in the setting $\Delta=0$, and will not consider the multiple testing problem (that is, we only test for an event at a single time $t$).
Proposition 1.
Fix any $T\geq 1$, any $\alpha\in(0,1)$, and any $t\in\{1,\cdots,T\}$. Assume $\sqrt{\alpha}T$ is an integer. Then there exist independent and stationary time series $\mathbf{X}$, $\mathbf{Y}\in\mathbb{R}^{T}$, and a function $\psi:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$, such that
$\mathbb{P}\left\{{p_{t}\leq\alpha}\right\}\geq\frac{\alpha}{\mathsf{mar}(t)}\cdot\left(\frac{1}{1+\sqrt{\frac{\alpha}{\mathsf{mar}(t)}}}\right)^{2}.$
Note that when $\alpha\ll\mathsf{mar}(t)$, this lower bound is approximately equal to $\frac{\alpha}{\mathsf{mar}(t)}$, which matches the upper bound shown in Theorem 1.
Proof.
Without loss of generality, we may assume $t\leq\frac{T+1}{2}$, so that $\mathsf{mar}(t)=(T/t)^{2}$. Let $m=\sqrt{\alpha}T$ and let $N$ be a large positive integer with $N>T$ (the value of $N$ will be specified below).
Define a time series $\mathbf{X}$ by taking the product of two components.
  • A cyclic component: let $I_{X}\sim\mathrm{Uniform}\left(\{0,\cdots,m+t-2\}\right)$. Define
    $\mathbf{X}^{\textnormal{cyclic}}=(X_{1}^{\textnormal{cyclic}},\cdots,X_{T}^{\textnormal{cyclic}}),\ \textnormal{ where }X_{i}^{\textnormal{cyclic}}={\mathbbm{1}}\left\{{\mathrm{mod}(I_{X}+i,m+t-1)\notin\{1,\cdots,t-1\}}\right\}.$
    This component is equivalent to sampling $T$ consecutive entries from the infinite periodic string
    $\cdots,\underbrace{1,\cdots,1}_{m\textnormal{ times}},\underbrace{0,\cdots,0}_{t-1\textnormal{ times}},\underbrace{1,\cdots,1}_{m\textnormal{ times}},\underbrace{0,\cdots,0}_{t-1\textnormal{ times}},\cdots$
  • A component that is monotonic with high probability: let $J_{X}\sim\mathrm{Uniform}\left(\{0,\cdots,N-1\}\right)$. Define
    $\mathbf{X}^{\textnormal{mono}}=(X_{1}^{\textnormal{mono}},\cdots,X_{T}^{\textnormal{mono}}),\ \textnormal{ where }X_{i}^{\textnormal{mono}}=N-\mathrm{mod}(J_{X}+i,N).$
    This component is equivalent to sampling $T$ consecutive entries from the infinite periodic string
    $\cdots,N,N-1,\cdots,2,1,N,N-1,\cdots,2,1,\cdots$
Let $\mathbf{X}=(X_{1}^{\textnormal{cyclic}}\cdot X_{1}^{\textnormal{mono}},\cdots,X_{T}^{\textnormal{cyclic}}\cdot X_{T}^{\textnormal{mono}})$, the elementwise product of the two components $\mathbf{X}^{\textnormal{cyclic}}$ and $\mathbf{X}^{\textnormal{mono}}$, with each of these two components generated independently. Since each component has a stationary distribution by construction, $\mathbf{X}$ is a stationary time series.
Next, let $\mathbf{Y}$ be generated independently from the same distribution as $\mathbf{X}$: we independently sample $I_{Y}\sim\mathrm{Uniform}\left(\{0,\cdots,m+t-2\}\right)$ and $J_{Y}\sim\mathrm{Uniform}\left(\{0,\cdots,N-1\}\right)$ independently, and define $\mathbf{Y}^{\textnormal{cyclic}}$, $\mathbf{Y}^{\textnormal{mono}}$, and $\mathbf{Y}$ analogously.
Let $\psi:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ be defined as $\psi(x,y)=(x+y)\cdot{\mathbbm{1}}_{{x>0}}\cdot{\mathbbm{1}}_{{y>0}}$. Consider the event that
$I_{X},I_{Y}\in\{0,\cdots,m-1\}\textnormal{ and }J_{X},J_{Y}\in\{0,\cdots,N-T-1\}.$ (13)
On this event, it holds that
$\mathrm{mod}(I_{X}+t,m+t-1)\in\{t,t+1,\dots,m+t-2,0\},$
which is the complement of $\{1,\cdots,t-1\}$. Thus, $X^{\textnormal{cyclic}}_{t}=1$. Analogously, $Y^{\textnormal{cyclic}}_{t}=1$. Meanwhile, since $N>T$, it holds that
$X_{i}^{\textnormal{mono}}=N-J_{X}-i,\textnormal{ and }Y_{j}^{\textnormal{mono}}=N-J_{Y}-j$
for all $i,j\in[T]$.
Therefore, on the event in (13), $\Psi_{t}=\psi(X_{t},Y_{t})=X_{t}+Y_{t}=2N-J_{X}-J_{Y}-2t>0$, and for each $i,j\in[T]$ we can calculate
  • If $X^{\textnormal{cyclic}}_{i}=0$ or $Y^{\textnormal{cyclic}}_{j}=0$, then $\Psi_{i,j}=0$, and so
    ${\mathbbm{1}}\left\{{\Psi_{i,j}\geq\Psi_{t}}\right\}={\mathbbm{1}}\left\{{0\geq\Psi_{t}}\right\}=0;$
  • If $X^{\textnormal{cyclic}}_{i}=Y^{\textnormal{cyclic}}_{j}=1$, then $\Psi_{i,j}=X^{\textnormal{mono}}_{i}+Y^{\textnormal{mono}}_{j}$, and so
    ${\mathbbm{1}}\left\{{\Psi_{i,j}\geq\Psi_{t}}\right\}={\mathbbm{1}}\left\{{N-J_{X}-i+N-J_{Y}-j\geq 2N-J_{X}-J_{Y}-2t}\right\}={\mathbbm{1}}\left\{{i+j\leq 2t}\right\}.$
Therefore, on the event in (13),
$p_{t}$ $=\frac{1}{T^{2}}\sum_{i,j=1}^{T}{\mathbbm{1}}\left\{{\Psi_{i,j}\geq\Psi_{t}}\right\}=\frac{1}{T^{2}}\sum_{i,j=1}^{T}{\mathbbm{1}}\left\{{X^{\textnormal{cyclic}}_{i}=Y^{\textnormal{cyclic}}_{j}=1,i+j\leq 2t}\right\}$
$\leq\frac{1}{T^{2}}\sum_{i,j=1}^{T}{\mathbbm{1}}\left\{{X^{\textnormal{cyclic}}_{i}=Y^{\textnormal{cyclic}}_{j}=1,i\leq 2t-1,j\leq 2t-1}\right\}$
$=\frac{1}{T^{2}}\left(\sum_{i=1}^{2t-1}{\mathbbm{1}}\left\{{X^{\textnormal{cyclic}}_{i}=1}\right\}\right)\left(\sum_{j=1}^{2t-1}{\mathbbm{1}}\left\{{Y^{\textnormal{cyclic}}_{j}=1}\right\}\right).$
Since $X^{\textnormal{cyclic}}_{t}=1$ by (13), by definition of $\mathbf{X}^{\textnormal{cyclic}}$ we can verify that $\sum_{i=1}^{2t-1}{\mathbbm{1}}\left\{{X^{\textnormal{cyclic}}_{i}=1}\right\}\leq m$, and the analogous bound holds for $\mathbf{Y}$ (see Figure 11 for an illustration of this calculation). Therefore, on the event defined in (13),
$p_{t}\leq\frac{m^{2}}{T^{2}}=\alpha.$
[…]
Figure 11: An illustration of ${\mathbbm{1}}\left\{{X^{\textnormal{cyclic}}_{i}=Y^{\textnormal{cyclic}}_{j}=1}\right\}$, over all integers $i,j$, with red shading indicating a value of $1$. Note that on the event in (13), it holds that $X^{\textnormal{cyclic}}_{t}=Y^{\textnormal{cyclic}}_{t}=1$ (i.e., the point $(t,t)$ lies in a $m\times m$ red shaded square, as in the illustration). On this event, no other red shaded squares overlap with the region $(i,j)\in\{1,\dots,2t-1\}\times\{1,\dots,2t-1\}$ (indicated by shading with blue lines).
Finally, by definition of the distributions of $I_{X},J_{X},I_{Y},J_{Y}$, the event in (13) holds with probability
$\left(\frac{m}{m+t-1}\right)^{2}\cdot\left(\frac{N-T}{N}\right)^{2}=\frac{\alpha T^{2}}{\left(t+\sqrt{\alpha}T-1\right)^{2}}\cdot\left(\frac{N-T}{N}\right)^{2}\\ =\frac{\alpha}{\left(\frac{t}{T}+\sqrt{\alpha}-\frac{1}{T}\right)^{2}}\cdot\left(1-\frac{T}{N}\right)^{2}=\frac{\alpha}{\mathsf{mar}(t)}\cdot\left(\frac{1-\frac{T}{N}}{1+\sqrt{\frac{\alpha}{\mathsf{mar}(t)}}-\frac{1}{t}}\right)^{2},$
and therefore,
$\mathbb{P}\left\{{p_{t}\leq\alpha}\right\}\geq\frac{\alpha}{\mathsf{mar}(t)}\cdot\left(\frac{1-\frac{T}{N}}{1+\sqrt{\frac{\alpha}{\mathsf{mar}(t)}}-\frac{1}{t}}\right)^{2}.$
Finally, by choosing a sufficiently large positive integer $N$, we can ensure that
$\frac{1-\frac{T}{N}}{1+\sqrt{\frac{\alpha}{\mathsf{mar}(t)}}-\frac{1}{t}}\geq\frac{1}{1+\sqrt{\frac{\alpha}{\mathsf{mar}(t)}}},$
which completes the proof. ∎